OpenStudy (anonymous):

Find the volume of the largest right circular cone that can be inscribed in a sphere of radius r.

OpenStudy (dumbcow):

 volume of cone is \[V = \frac{\pi}{3} x^2 h\] h = r+y and x^2 + y^2 = r^2 or x^2 = r^2 - y^2 \[\rightarrow V = \frac{\pi}{3} (r^2 -y^2)(r+y)\] maximize volume by differentiating with respect to "y" and setting equal to 0